7

7. Find the volume of the solid under the paraboloid 2 = 4-12 - 3y over the square R= [0, 1] x [0,1).
Answer #1

Giren 2=4-xe - 3y², R=[0, 1] x [..] &(4-v2 dy dx v=SS z dv 4-12-344) 6 44-x²y- 4(1-0)-x{1-03-[1-0] da 6 (4 - x²-1) de 7*-